OpenStudy (anonymous):

Solve. 2x2 = 8 quadratics

OpenStudy (anonymous):

2x^2=8? div by 2 x^2=4 take sqrt x=sqrt(4)

OpenStudy (anonymous):

Don't forget \[\pm\]

OpenStudy (anonymous):

sqrt(x) implies that right?

OpenStudy (anonymous):

but yea dont forget it.

OpenStudy (anonymous):

2x^2 = 8 x^2 = 8/2 x^2 = 4 x = Sqrt 4 x = ±2